How many gallons of 70% antifreeze should be mixed with 10 gallons of 30% antifreeze to obtain a 45% antifreeze mixture?
Here's what I have
(0.30)(10gal)+(0.70)(x)=(0.45)(10+x)
After I solve this I get 6 gallons but when I check the solution I don't get 45% so this cannot be the right way to solve this problem. Please help.

Answer by richwmiller(17219)

We have 10 gal of 30% antifreeze
We want to add some 70% antifreeze to make 45% antifreeze
0.7*x+0.3*10=0.45(10+x)
0.7*x+3=4.5+0.45x
0.7x-0.45x=4.5-3
0.25x=1.5
x=6 gal at 70% (4.2) is added to 10 gal at 30%(3) to make 45% of 16 gal(7.2)
check
Remember it is 45% of 16 gallons that is .45*16 which is 7.2
0.7*6+0.3*10=0.45(10+6)
4.2+3=0.45(16)
7.2=7.2
ok